Let y = -2cos(3(x-pi/2)).
What is the amplitude??
What is the period??
For the phase shift, which way is the graph shifted?
How far is it shifted?
(Note: you can enter pi in the answer if applicable)

Answer by stanbon(48558)

Let y = -2cos(3(x-pi/2))
y = -2cos(3x-(3/2)pi)
What is the amplitude??
Ans: |-2|=2
------------------------------
What is the period??
Ans: P = 2pi/b = 2pi/3 = (2/3)pi
--------------------------------------------
For the phase shift, which way is the graph shifted?
To the right.
-------------------------
How far is it shifted?
-c/a = (-3/2)/3 = 1/2 radian to the right
------------------------------
Cheers,
Stan H.